# Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach

### v3

@@ Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a protocol that enables exponential acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the quantum Mpemba effect. Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol eliminates multiple relaxation modes simultaneously. Specifically, using a simple depolarizing channel, we show that the relaxation of a dissipative transverse-field Ising chain can be significantly accelerated, underscoring the experimental feasibility of the protocol. This framework provides a versatile tool for controlling quantum relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
@@
Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies. The timescales of such relaxations often determine the feasibility and performance of quantum devices. For example, in quantum thermodynamics, the power output of heat engines or transport junctions is limited by how rapidly the working substance reaches its steady state in each cycle [1, 2, 3]. In addition, fast relaxation facilitates ground state laser cooling [4, 5, 6] and the reliable preparation of quantum states [7]. Faster relaxation can also lead to more efficient quantum algorithms [8], particularly in dissipative computational architectures where the output is encoded in the system’s stationary state [8, 9]. Slow relaxation inevitably allows other unwanted dissipative dynamics to consume substantial time resources, compromising the efficiency of the final stabilized state [10, 11]. In contrast, slow relaxation is sometimes desirable—such as when the computational task targets metastable states [12, 13, 14, 15], whose prolonged lifetimes are essential for robust information storage or approximate optimization.
@@
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [16] and nonequilibrium steady states [17, 18, 11]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [19], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] and quantum systems [10, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42], see also [43, 44] for recent reviews. However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real. Moreover, careful initial-state design is a delicate task, since it relies on detailed knowledge of the original initial state and the system dynamics, which is typically not known a prior in complex quantum systems. These issues limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
@@
In this study, we address these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [45, 46, 47]. Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [48, 49, 9, 50], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [51, 52, 53, 54, 55, 56, 57, 58]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this protocol can exponentially accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous elimination of multiple relaxation modes—a capability absent in previous approaches. Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
@@
Setup.—We consider a general open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics in the interaction picture are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
@@
Here, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment. The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$. Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization
@@
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}c_{k}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $||\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right),$ so the characteristic relaxation timescale is given by $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. For clarity of exposition, we first consider the simple case where the second eigenvalue $\lambda_{2}$ is real and non-degenerate, as in [10]. Later, we will generalize the discussion to include cases where the subleading eigenvalues form complex conjugate pairs.
Reset protocol to accelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that suppresses its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$. The modified dynamics during the reset phase is governed by a new Lindbladian $\mathcal{L}_{r}$ [45], defined as
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho.$ (4)
Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i}^{r,\alpha}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter). The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. Consequently, the modified Lindbladian acts as
@@
This induces a uniform spectral shift by $-r$ for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged. In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
@@
where $d_{k}=\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$. By properly choosing an optimal $t^{*}_{s}$, we can set $c_{2}^{r}(t^{*}_{s})=0$, thereby eliminating the overlap with the slowest decaying mode. Solving $c_{2}^{r}(t_{s})=0$ yields
$t^{*}_{s}(r)=\frac{1}{r-\lambda_{2}}\ln\left[1-\frac{c_{2}(r-\lambda_{2})}{r\cdot d_{2}}\right].$ (8)
By an appropriate choice of $\rho_{\delta}$ such that $c_{2}/d_{2}\leq 0$, $t^{*}_{s}(r)\geq 0$ can be assured. After the reset phase ($t>t_{s}$), the system returns to original Lindbladian evolution, now without the contribution from the slowest mode. The full evolution reads:
$\rho^{r}(t)=$ $\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t)R_{k}$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s})R_{k},$ (9)
where $\Theta(x)$ is the Heavside step function. This protocol with an optimal $t_{s}=t^{*}_{s}$ reduces the relaxation time to
$\tau_{{\rm rel}}\sim t^{*}_{s}+\frac{1}{|{\rm Re}(\lambda_{3})|},$ (10)
which is generically shorter than the unperturbed timescale $\tau_{{\rm{\rm rel}}}\sim\frac{1}{|\text{Re}(\lambda_{2})|}$. Eq. (9) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
Our protocol also enables controlled deceleration of relaxation. By choosing $\rho_{\delta}$ and $t_{s}$ to amplify the magnitude of $c_{k}^{r}(t)$, one can prolong the time during which the system remains away from the steady state. This capability could be useful in quantum metrology and computation, where long-lived metastable states play a functional role. Our reset protocol thus provides a general framework for controlling relaxation processes in open quantum systems.
Extension to complex-conjugate relaxation modes.—We now consider the more general case where the dominant relaxation modes form a complex-conjugate pair. In this case, the long-time dynamics of the density matrix takes the form
$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{\dagger}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
@@
where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate. Most known strategies are unable to simultaneously suppress both components of such complex relaxation modes.
Our reset protocol, however, naturally addresses this limitation. The modified coefficients under quantum reset are given by
$c_{2}^{r}(t)=\left[c_{2}-\frac{r\cdot d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\cdot d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t}$ , (12a)
$c_{2}^{r,*}(t)=\left[c_{2}^{*}-\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\cdot d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}$ (12b)
Importantly, we observe that these coefficients remain complex conjugates of each other for all $t$, i.e., $[c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$. Therefore, eliminating the entire complex mode reduces to a single complex condition:
$c_{2}^{r}(t_{s})=0\Leftrightarrow\text{Re}[c_{2}^{r}(t_{s})]=\text{Im}[c_{2}^{r}(t_{s})]=0,$ (13)
because $c_{2}^{r}(t_{s})=0\Rightarrow c_{2}^{r,*}(t_{s})=[c_{2}^{r}(t_{s})]^{*}=0$. In some cases, this condition can be exactly satisfied by tuning $r$ and $t_{s}$. Although Eq. (13) does not always admit an exact solution, in practice, it suffices to reduce $|c_{2}^{r}(t_{s})|$ to an exponentially small value ($\text{Re}[c_{2}^{r}(t_{s})]\ll 1,\ \text{Im}[c_{2}^{r}(t_{s})]\ll 1$), which still leads to exponential acceleration of relaxation. Additionally, we show in [59] that, if the condition
${\rm Re}(c_{2}^{*}d_{2})\leq|c_{2}|^{2}$ (14)
is satisfied, $|c_{2}^{r}(t_{s})|<|c_{2}|$ can always be achieved for any positive $t_{s}$ that below a certain threshold. This is a weak condition that be readily satisfied for many choices of $\rho_{\delta}$. It is strictly weaker than the commonly satisfied condition $c_{2}/d_{2}\leq 0$, which, in the case where $\lambda_{2}$ is real, guarantees the existence of a $t_{s}^{*}$ such that $c^{r}_{2}(t_{s}^{*})=0$. Eq. (14) thus provides a sufficient condition for reset-induced Mpemba-type acceleration [20]. It implies that fine-tuning of either $\rho_{\delta}$ or $t_{s}$ is unnecessary for achieving acceleration. We illustrate this behavior concretely in the following examples.
@@
Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. While minimal, this example is crucial in many practical settings, as qubit reset often constitutes the bottleneck in large-scale quantum processes [60, 58]. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads $H=E|1\rangle\langle 1|+\Omega(\sigma^{+}+\sigma^{-})$, where $\Omega$ is the intrinsic coherent coupling between the two levels, $\sigma^{+}=|1\rangle\langle 0|$ and $\sigma^{-}=|0\rangle\langle 1|$. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$. Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or converse). Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
@@
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. In the absence of intrinsic transitions ($\Omega=0$), the Hamiltonian is diagonal and the relaxation process reduces to pure decoherence. In this case, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59]. When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state to ensure $t_{s}(r)\geq 0$, the two slowest decaying modes form a complex-conjugate pair:
@@
This overlap cannot be entirely eliminated in finite time, but it decays exponentially fast, leading to a relaxation timescale
$\tau_{{\rm rel}}\sim\frac{1}{r}+\frac{1}{|\text{Re}(\lambda_{3})|},$ (16)
where $1/r$ characterizes the reset protocol timescale, and $1/|\text{Re}(\lambda_{3})|$ represents the accelerated relaxation timescale with the slowest mode eliminated. Clearly, $\tau_{{\rm rel}}<1/|{\rm Re}(\lambda_{2})|$ for sufficiently large $r$. By appropriately tuning the reset rate, the relaxation timescale can approach $1/|\rm Re(\lambda_{3})|$ arbitrarily closely. Using Eq. (9), the full dynamics under our reset protocol can thus be determined.
Since the condition $c_{2}^{r}(t_{s})=0$ cannot be satisfied exactly in this case—though $|c_{2}^{r}(t_{s})|$ can be made exponentially small by choosing $t_{s}\sim 1/r$—we instead define $t_{s}$ such that a faster decaying mode is suppressed. Specifically, solving $c_{4}^{r}(t_{s})=0$ gives
$t_{s}=\frac{\ln\left[1+\frac{\gamma_{1}k_{\text{tot}}}{r}\left(\frac{1}{\gamma_{0}+\gamma_{1}e^{(\beta_{0}-\beta_{\text{env}})E}}-\frac{1}{\gamma_{0}+\gamma_{1}}\right)\right]}{k_{\text{tot}}},$ (17)

### v4

@@ Abstract
Slow relaxation processes spanning widely separated timescales pose fundamental challenges for probing steady-state properties and engineering functional quantum systems, such as quantum heat engines and quantum computing devices. We introduce a protocol that enables significant acceleration of relaxation in general Markovian open quantum systems by temporarily coupling the system to a reset channel, inspired by the Mpemba effect. Crucially, this acceleration persists even when the slowest decaying Lindbladian modes form complex-conjugate pairs. Unlike previous approaches, which typically target a single mode, our protocol may suppress multiple relaxation modes simultaneously. This framework provides a versatile and experimentally feasible tool for controlling relaxation timescales, with broad implications for quantum thermodynamics, computation, and state preparation.
@@
Introduction.—Relaxation processes in open quantum systems, wherein a driven or interacting system settles into an equilibrium or a nonequilibrium steady state, are of fundamental importance in nonequilibrium physics and for practical quantum technologies. The timescales of such relaxations often determine the feasibility and performance of quantum devices. For example, in quantum thermodynamics, the power output of cyclic heat engines or transport junctions can be limited by how rapidly the working substance relaxes to its steady state in each cycle [1, 2]. In addition, fast relaxation facilitates ground state laser cooling [3, 4, 5] and the reliable preparation of quantum states [6]. Faster relaxation can also lead to more efficient quantum algorithms [7], particularly in dissipative computational architectures where the output is encoded in the system’s stationary state [7, 8]. Slow relaxation inevitably allows other unwanted dissipative dynamics to consume substantial time resources, compromising the efficiency of the final stabilized state [9, 10]. In contrast, slow relaxation is sometimes desirable—such as when the computational task targets metastable states [11, 12, 13, 14], whose prolonged lifetimes are essential for robust information storage or approximate optimization.
@@
Researchers have developed various techniques to accelerate relaxation toward both thermal equilibrium [15] and nonequilibrium steady states [16, 17, 10]. Most of these methods, however, are restricted to specific systems and rely on intricate external controls. For instance, some approaches require auxiliary Hamiltonians or full knowledge of the system Hamiltonian, limiting their applicability to simple setups. A more general and elegant strategy is to accelerate relaxation by tailoring the initial state distribution, without the need for continuous external driving. A particularly intriguing example of this approach is inspired by the so-called Mpemba effect [18], a counterintuitive relaxation anomaly in which a hotter system can cool faster than a colder one. This phenomenon implies the existence of an optimal initial state that minimizes the relaxation timescale. Theoretical frameworks for identifying such optimal states have been proposed in both classical [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and quantum systems [9, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], see also [42, 43] for recent reviews. However, existing quantum implementations typically impose strict constraints, such as assuming pure initial states or requiring the second-largest eigenvalue of the Lindbladian to be real. Moreover, careful initial-state design is delicate, since it relies on fine-tuning of control parameters and detailed knowledge of the initial state and system dynamics, which are typically not known a priori in complex quantum systems. These issues limit their applicability in general settings. Thus, a general and practical method for inducing faster relaxation in open quantum systems, which is robust with respect to different initial states, remains elusive.
@@
In this study, we contribute to addressing these challenges by proposing an experimentally feasible protocol that can both accelerate and decelerate general quantum relaxation processes from any initial state, by resetting the system to a specified (possibly mixed) target state [44, 45, 46]. Here, this type of quantum reset operation can be realized through engineered dissipation, i.e., quantum reservoir engineering [47, 48, 8, 49], and should be distinguished from the conventional reset typically used for qubit initialization in quantum computation [50, 51, 52, 53, 54, 55, 56, 57]. In our protocol, we introduce a finite-duration reset phase into the open-system dynamics by temporarily coupling the system to a reset channel. We show that this protocol can significantly accelerate relaxation, even when starting from mixed initial states and in the presence of complex decay modes. Remarkably, our method allows for the simultaneous suppression of multiple relaxation modes—a capability absent in previous approaches. Moreover, selective application of the protocol enables the system of interest to remain in metastable states for extended periods, offering a systematic route to suppress relaxation.
@@
Setup.—We consider a general Markovian open quantum system defined on a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) master equation $d\rho/dt=\mathcal{L}(\rho),$ where the Lindbladian superoperator $\mathcal{L}$ is given by
@@
Here, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment. The evolution can be analyzed via the spectral decomposition of $\mathcal{L}$, whose eigenvalues are denoted $\lambda_{k}$ and ordered such that $0\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1})$. Assuming that $\mathcal{L}$ is diagonalizable, let $R_{k}$ and $L_{k}$ denote the corresponding right and left eigenmatrices, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{\star}L_{k},\ k=1,...,d^{2}$, with biorthogonal normalization [58]
@@
$\rho(t)=e^{t\mathcal{L}}[\rho_{0}]=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}R_{k},$ (3)
where $c_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{0})$. The unique stationary state of the open quantum system $\rho_{\rm ss}$ is given by $\rho_{\rm ss}=\lim_{t\rightarrow\infty}\rho(t)=R_{1}$, assuming $\lambda_{1}=0$ is non-degenerate. In the long-time limit, the relaxation is dominated by the slowest decaying mode, and the deviation from stationarity obeys $\rho(t)-\rho_{\rm ss}||\sim\exp\left(-|\text{Re}\lambda_{2}|t\right)$.
Reset protocol to accelerate or decelerate relaxation processes.—Recognizing that the slowest decaying mode governs the relaxation timescale, we propose a quantum reset protocol that can suppress or promote its excitation from a general initial state via a finite-duration reset phase, thereby enabling exponential acceleration of relaxation. Specifically, we let the system evolve under modified dynamics for a finite time $t_{s}$, during which it is stochastically reset to a chosen state $\rho_{\delta}$ at random times, with events following a Poisson process of rate $r$. After time $t_{s}$, the reset channel is turned off, and the system resumes evolving under the original Lindbladian $\mathcal{L}$. That is, the system density matrix $\rho(t)$ evolves under the reset channel for $t\in[0,t_{s}]$ and decouples from the channel for $t>t_{s}$. The modified dynamics during the reset phase $t\in[0,t_{s}]$ is governed by a new Lindbladian $\mathcal{L}_{r}$ [44], defined as
$\mathcal{L}_{r}(\rho):=\mathcal{L}(\rho)+\mathcal{D}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}(\rho)\rho_{\delta}-r\rho,$ (4)
where $\mathcal{D}_{r}(\rho)=r[\text{Tr}(\rho)\,\rho_{\delta}-\rho]$. Expressing $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\langle\psi_{\alpha}|$, $\mathcal{L}_{r}$ is equivalent to introducing jump operators $J_{i,\alpha}^{r}=\sqrt{rp_{\alpha}}|\psi_{\alpha}\rangle\langle\phi_{i}|$ into $\mathcal{L}$, where $|\phi_{i}\rangle$ form an auxiliary complete orthonormal basis (see End Matter). The term $r\text{Tr}(\rho)\rho_{\delta}$ contributes only to the stationary mode and vanishes on all traceless eigenmodes $R_{i}$ with $i\geq 2$, which satisfy $\mathrm{Tr}(R_{i})=0$ due to $\mathrm{Tr}(L_{1}^{\dagger}R_{i})=0$ and $L_{1}^{\dagger}=\mathbb{I}$ [cf. (2)]. Consequently, the modified Lindbladian acts as
@@
This induces a uniform spectral shift by r-r for all non-stationary eigenmodes, leaving the eigenmatrices ($R_{i}$, $i\geq 2$) themselves unchanged. In contrast, the stationary mode $R_{1}$ retains eigenvalue zero but is modified under reset. Let $\rho^{r}(t):=e^{t\mathcal{L}_{r}}\rho_{0}$ be the state at time $t$ under reset dynamics. Its spectral decomposition reads:
@@
where $d_{k}\equiv\text{Tr}(L_{k}^{\dagger}\rho_{\delta})$ is the overlap coefficient of $\rho_{\delta}$. Notably, Eq. (Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach) can be generalized to the dynamics of observables [59], which may be more convenient to measure.
After the reset phase ($t>t_{s}$), the system returns to the original Lindbladian evolution. The dynamics for $t>t_{s}$ then reads:
$\rho^{r}(t)=\rho_{\rm ss}+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}R_{k}.\quad(t>t_{s})$ (8)
Importantly, the modified spectral decomposition, Eq. (8), takes the same form as the original one, Eq. (3), which allows us to directly compare the relaxation dynamics with and without reset. From this comparison, it is clear that for $t>t_{s}$ our protocol is equivalent to the original Lindbladian evolution, but with different overlap coefficients,
$\rho(t)=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}c^{\prime}_{k}e^{\lambda_{k}t}R_{k}.\quad(t>t_{s})$ (9)
Here, $c_{k}^{\prime}\equiv c^{r}_{k}(t_{s})$ are determined by the reset protocol.
Thus, our protocol controls relaxation dynamics by tuning the overlap coefficients $c_{k}$, without requiring any special preparation of the initial state. By choosing an appropriate $\rho_{\delta}$ and varying $r$ and $t_{s}$, one can tune the amplitudes of multiple relaxation modes. In particular, ensuring $|c_{2}^{r}(t_{s})|<|c_{2}|$ ($|c_{2}|$) suppresses (enhances) the dominant mode, which is sufficient for acceleration (deceleration), ultimately bringing the system closer to (farther from) stationarity [19, 32]. Specifically, the dominant mode is fully eliminated and exponential speed-up occurs when $|c_{2}^{r}(t_{s})|=0$; this, however, requires fine-tuning and prior knowledge and is therefore challenging to realize in large systems. We thus focus on the weaker but practically relevant scenario $|c_{2}^{r}(t_{s})|<|c_{2}|$.
Sufficient conditions for suppression or promotion of relaxation modes.—We derive a sufficient condition [59],
${\rm Re}(c_{2}^{*}d_{2})<|c_{2}|^{2},$ (10)
under which there always exists a threshold $t_{c}>0$ such that, for any $r>0$ and any $t_{s}\in(0,t_{c}]$, one has $|c_{2}^{r}(t_{s})|<|c_{2}|$. That is, the relaxation is accelerated for all $t_{s}$ in this range. If the condition is violated, there exists another threshold $t_{c}^{\prime}>0$ with $|c_{2}^{r}(t_{s})|\geq|c_{2}|$ for $t_{s}\in(0,t_{c}^{\prime}]$, leading to deceleration. However, Eq. (10) is not necessary for acceleration, since one may still have $|c_{2}^{r}(t_{s})|<|c_{2}|$ for $t_{s}>t_{c}^{\prime}$, when it is not satisfied.
Several remarks are in order. First, the condition applies both when $\lambda_{2}$ is real and when it forms a complex-conjugate pair. In the latter case, the long-time dynamics takes the form
$\rho(t)=\rho_{\rm ss}+e^{\lambda_{2}t}c_{2}R_{2}+e^{\lambda_{2}^{*}t}c_{2}^{*}R_{2}^{\dagger}+\mathcal{O}(e^{\lambda_{3}t}),$ (11)
@@
where $c_{2}=\text{Tr}(L_{2}^{\dagger}\rho_{0})$ and $c_{2}^{*}=\text{Tr}(L_{2}\rho_{0})$ is its complex conjugate. Most existing strategies cannot simultaneously suppress both components of such complex relaxation modes. Our protocol, however, naturally addresses this limitation. The modified coefficients under reset are
$c_{2}^{r}(t)$ $=\left[c_{2}-\frac{r\,d_{2}}{r-\lambda_{2}}\right]e^{-rt}+\frac{r\,d_{2}}{r-\lambda_{2}}e^{-\lambda_{2}t},$ (12a)
$c_{2}^{r,*}(t)$ $=\left[c_{2}^{*}-\frac{r\,d_{2}^{*}}{r-\lambda_{2}^{*}}\right]e^{-rt}+\frac{r\,d_{2}^{*}}{r-\lambda_{2}^{*}}e^{-\lambda_{2}^{*}t}.$ (12b)
These coefficients remain complex conjugates for all $t$, i.e., $c_{2}^{r}(t)]^{*}=c_{2}^{r,*}(t)$. Eq. (10) then ensures that $|c_{2}^{r,*}(t_{s})|=|c_{2}^{r}(t_{s})|\leq|c_{2}|=|c_{2}^{*}|$ for a moderate $t_{s}$.
Second, Eq. (10) is a weak condition satisfied by many choices of $\rho_{\delta}$. In particular, it does not require that $\rho_{\delta}$ be closer to the steady state than $\rho_{0}$, nor that the overlap coefficient $|d_{2}|$ be smaller than $|c_{2}|$. For example, when $\lambda_{2}$ is real, if ${\rm Re}(d_{2})$ and $c_{2}$ have opposite signs, the condition is satisfied even when $|d_{2}|$ is arbitrarily large. In Example 2 we show that the condition is typically satisfied for a large fraction of randomly chosen initial states.
Third, neither $r$ nor $t_{s}$ needs to be precisely tuned in advance to achieve acceleration or deceleration. Provided that Eq. (10) holds, any positive $r$ and moderately small $t_{s}$ are sufficient for acceleration, and $t_{s}$ can be further optimized a posteriori based on the observed dynamics. Our protocol therefore provides a general and robust framework for controlling relaxation in open quantum systems, without the need for fine-tuning.
Finally, it is straightforward to see that
${\rm Re}(c_{k}^{*}d_{k})<|c_{k}|^{2}$ (13)
is a sufficient condition for suppression of the $k$-th mode. Since multiple such conditions for different $k$ can be satisfied simultaneously, our protocol allows simultaneous suppression (or promotion) of multiple relaxation modes. In [59] we provide numerical evidence that multiple conditions can indeed be satisfied in parallel, using the model and protocol of Example 2.
So far we have focused on practical scenarios where fine-tuning is not possible and neither the system dynamics nor the initial state are known a priori. If these constraints are lifted, as in most previous studies, our protocol can fully eliminate the dominant relaxation mode ($|c_{2}^{r}(t_{s})|=0$) by choosing an optimal $t_{s}=t_{s}^{*}$. When $\lambda_{2}$ is real, the exact expression is [59, 24]
$t^{*}_{s}=\frac{1}{r-\lambda_{2}}\ln\!\left(1-\frac{c_{2}(r-\lambda_{2})}{rd_{2}}\right).$ (14)
A sufficient condition ensuring that $t^{*}_{s}\geq 0$ is $c_{2}/d_{2}\leq 0$. Further details, including the case of complex-conjugate pairs, are given in [59].
@@
Example 1.—To illustrate our protocol, we first consider a minimal example: a two-level quantum system, which may represent a qubit or a spin-$1/2$ particle. While minimal, this example is crucial in many practical settings, as qubit reset often constitutes the bottleneck in large-scale quantum processes [60, 57]. We set the ground state energy $E_{0}=0$ and excited state energy $E_{1}=E$, such that the system Hamiltonian reads $H=E|1\rangle\langle 1|+\Omega(\sigma^{+}+\sigma^{-})$, where $\Omega$ is the intrinsic coherent coupling between the two levels, $\sigma^{+}=|1\rangle\langle 0|$ and $\sigma^{-}=|0\rangle\langle 1|$. The system is coupled to a thermal environment through the Lindblad jump operators $J_{0}=\sqrt{\gamma_{0}}\sigma^{+}$ and $J_{1}=\sqrt{\gamma_{1}}\sigma_{-}$. Here, $\gamma_{0}$ ($\gamma_{1}$) is the transition rate from state $|0\rangle$ to state $|1\rangle$ (or vice versa). Assuming the rates satisfy detailed balance with inverse temperature $\gamma_{0}e^{-\beta_{\text{env}}E_{0}}=\gamma_{1}e^{-\beta_{\text{env}}E_{1}}$, i.e., $\gamma_{0}=\gamma_{1}e^{-\beta_{\text{env}}E},$ the system evolves to a unique equilibrium state $\rho_{\rm eq}=\frac{e^{-\beta_{\text{env}}H}}{\text{Tr}(e^{-\beta_{\text{env}}H})}$ irrespective of the initial condition. We initialize the system in a general mixed state:
@@
where the off-diagonal terms $ke^{\pm i\phi}$ quantify initial quantum coherence. When $\Omega=0$, all eigenmodes and eigenvalues of the Lindbladian can be computed analytically [59]. When we apply our reset protocol with the reset state $\rho_{\delta}=|0\rangle\langle 0|$, corresponding to the ground state, the two slowest decaying modes form a complex-conjugate pair:
@@
These overlaps are exponentially suppressed compared with the original overlap, i.e. $c_{2}^{r}(t)=e^{-rt}c_{2}$ and $c_{2}^{r,*}(t)=e^{-rt}c_{2}^{*}$. Clearly, the relaxation is accelerated for any $r$ and $t_{s}$.